Question: The cost of the telephone call for 8 minutes between two countries is Rs.315 and a 10 minutes call costs rupees 385. Find the cost of first 3 minutes and the rate of each additional minute call.
Answer by edjones(8007)

385-315=Rs 70 for 2 min.
Each additional min is 70/2=35
35*5=175 cost of 5 additional min.
315-175=Rs 140 cost of the 1st 3 min.
.
Ed